write the explicit formula for the sequence. Then find the fifth term in the sequence.
asub1=3, r=-3

Answer by Edwin McCravy(20081)

an = a1·rn-1

a5 = a1·(-3)5-1

a5 = 3·(-3)4

a5 = 3·81 = 243

Check it by writing them all out, multiplying by -3 each time:

First term  =              +3
Second term =  +3 × -3 =   -9
Third term  =  -9 × -3 =  +27
Fourth term = +27 × -3 =  -81
Fifth term  = -81 × -3 = +243

So 243 is correct.

Edwin